# PaintScene4D: Consistent 4D Scene Generation from Text Prompts

### v1

@@ 4.3 Scene Completion with Temporal Consistency
Figure 3: Gallery of Results. PaintScene4D successfully generates 4D scenes that maintain both view consistency and temporal coherence. Here, we present visualizations of each scene from 2 fixed viewpoints across varying timestamps. The horizontal axis represents the time progression, while the vertical axis represents different viewpoints. Due to space constraints, we display results for only four scenes. Additional results are provided in the supplementary materials for further reference.
Figure 4: Text-to-4D Comparisons. We compare our approach with state-of-the-art text-to-4D generation methods. While both baseline methods produce scenes that broadly align with the text prompts, they lack essential fine details. Specifically, 4D-fy shows minimal motion and limited detail, whereas Dream-in-4D captures dynamics more effectively but produces stylized, cartoon-like renderings, which are less suitable for realistic applications. In contrast, our method synthesizes photorealistic 4D scenes that faithfully follow the input text prompt while presenting significant, realistic dynamics within the scene.
Figure 5: Camera Control. PaintScene4D demonstrates strong explicit camera control capabilities. Input text prompt: “A rabbit flying a drone in the middle of mountains.” In contrast, to guide the T2V model, we append camera motion directives such as “The camera tilts to the right” or “The camera moves upwards” to the text prompt. However, due to implicit handling of camera motion, T2V often fails to generate precise or controllable camera movements. Our approach, once trained, allows for flexible camera trajectories within the bounds of the input training cameras, achieving precise and repeatable control over camera movements for the same scene.
@@ 4.4 Training and Optimization
After performing all warping and inpainting operations across views and timestamps, we establish a comprehensive camera network, where each camera contains video frames as captured from its respective viewpoint. Importantly, this multi-view setup is constructed without the need for model-specific training. Using this multi-view spatial information and temporal dynamics, we employ a 4D rendering approach to synthesize novel perspectives of the scene. For the rendering, we leverage the 4D Gaussians framework [51], which represents the 4D space using a deformable network. The renderer takes Gaussian parameters, along with the timestamp, and compute the timestamp-conditioned deformation of these parameters. This approach enables continuous modeling of deformation, facilitating smooth interpolation between timestamps during novel view synthesis. At test time, any desired viewpoint and timestamp can be selected to generate a novel view.
@@ 5 Experiments
5.1 Implementation Details
Our optimization framework comprises two stages: initially reconstructing a network of cameras, each associated with its respective timeframe view, followed by training a 4D renderer. Specifically, we construct a network of 25 cameras and utilize videos spanning 50 timestamps. All experiments were conducted on a single A100 GPU. The complete warping and inpainting process, which operates without any additional training, requires approximately two hours. Following this, the 4D renderer is trained in about one hour, resulting in a total of approximately 3 hours to complete the training and generate novel views along any desired trajectory. This duration is significantly shorter than the training time required by recent state-of-the-art methods: Dream-in-4D requires over four hours, while 4Dfy takes over 20 hours, despite producing only object-level 4D renderings. To initialise the scene and establish motion priors for 4D reconstruction, we utilize CogVideoX-5b [54]. For depth estimation, DepthCrafter [16] is employed, as it produces consistent depth estimates across video frames, enabling reliable warping. Perspective Fields [20] is used to estimate the camera intrinsics for the generated video. For the segmentation model to distinguish foreground and background, we use GroundingSAM-2[37].
@@ 5.2 Baselines and Evaluation Metrics
In the absence of open-source implementations for text-to-4D scene-level generation, we benchmark our approach against state-of-the-art text-to-4D object-level generation methods, namely 4Dfy [2] and Dream-in-4D [62], across a varied set of 20 prompts. To assess the effectiveness of our proposed approach, we utilize the CLIP Score [35] alongside a structured user study.
CLIP Score.   The CLIP Score [32] assesses alignment between an input prompt and generated visual content by calculating the cosine similarity between CLIP’s text and visual embeddings [35]. Scored between 0 and 100, a higher value indicates a closer match. We compute the CLIP Score for 4Dfy, Dream-in-4D, and our method by rendering videos from each prompt, evaluating each frame with CLIP ViT-B/32, and averaging scores across all frames and prompts for consistency.
User Study.   A comprehensive user study was conducted through Google Forms, involving 30 evaluators per video pair. Each evaluator was presented with three anonymized videos, each capturing a dynamic scene from a camera moving along a circular trajectory. The videos were accompanied by the original text prompt. Evaluators were shown the renderings from 4Dfy, Dream-in-4D, and our approach simultaneously, and were asked to rate their preferences based on 4 criteria: motion realism, video-text alignment, high dynamicity, general realism. Detailed results are presented in Table 1.
Table 1: Quantitative results. We compare our method against 4D-fy and Dream-in-4D. The methods are evaluated in terms of CLIP Score (CLIP) and human preference based on motion realism(MR), video-text alignment (VTA), High Dynamicity (HR), general realism (GR), and overall preference (Overall). The reported human preference is the percentage of users who voted for the respective method in a head-to-head comparisons.
@@
Method CLIP MR VTA HR GR Overall
4D-fy [2] 31.82 2% 11% 5% 7% 7%
Dream-in-4D [62] 28.14 13% 14% 17% 2% 11%
Ours 36.02 85% 75% 78% 91% 82%
5.3 Text-to-4D Generation
In Figure 4, we visualize spatio-temporal renderings produced by our method compared to 4D-fy and Dream-in-4D. Although all approaches are capable of synthesizing 4D scenes, 4D-fy and Dream-in-4D focus on object-level renderings and lack fine spatial details. Our approach, by contrast, generates scene-level 4D reconstructions in a significantly reduced time, producing realistic renderings. Notably, 4D-fy struggles to model realistic motion, while Dream-in-4D produces cartoonish effects that diminish realism. In contrast, our method achieves high photorealistic quality across both spatial and temporal dimensions. We also a present a gallery of our results in Figure 3. Quantitative metrics indicating both CLIP score and User study are reported in Table 1. Our method outperforms both 4D-fy and Dream-in-4D in terms of CLIP Score and user study preferences. Evaluators showed a statistically significant preference for PaintScene4D due to its higher motion realism, photorealistic rendering of both foreground and background, overall realism, and better video-text alignment.
@@ 5.4 Explicit Camera Control
To assess camera control, we compared our framework with other text-to-video (T2V) models as illustrated in Figure 5. We input the same text prompts into the T2V model twice, adjusting only the camera movement description to direct it to “tilt towards the right” in one case and “move upwards” in the other. This setup allows us to examine the models’ ability to interpret and execute nuanced camera control instructions. Our observations reveal two key limitations of T2V models. First, even with a fixed seed, the T2V model generates different scenes for each altered prompt. Secondly, although the model simulates an upward camera movement in the second case, it lacks explicit control over the degree of camera motion. In contrast, our approach enables explicit, consistent control over camera trajectory within the same scene and motion dynamics, leveraging 4D modeling for precise camera manipulation.
Figure 6: Ablation Study. We demonstrating that each of our proposed elements is essential for mitigating artifacts and inconsistencies, resulting in smooth consistent renderings.
@@ 6 Ablation Study
We present an ablation study to analyze the components of our PaintScene4D framework, with the results shown in Figure 6.
Depth Alignment Module:   The inclusion of the depth alignment module is crucial for maintaining the geometric consistency of the foreground. During the warping process, all frames are utilized, and any depth inconsistencies across frames result in error accumulation, leading to noticeable artifacts, particularly at the foreground boundaries.
Farthest View Sampling:   In PaintScene4D, we select the farthest view at each step of the warping process to maximize the area inpainted. Omitting this step causes severe degradation near the edges of the foreground, such as the panda’s boundary, where needle-like artifacts emerge due to the Gaussian splatting process.
Consistent Inpainting Module:   Temporal consistency in inpainting is essential for coherent 4D scene generation. Without this module, inpainting becomes inconsistent at the boundaries of objects (e.g., the panda) across different timestamps, leading to significantly degraded renderings.
@@ 7 Conclusion
We introduce PaintScene4D, a novel training-free framework for generating photorealistic 4D scenes from a single text prompt. Our method addresses the challenges of spatial and temporal inconsistencies and enables the generation of novel views along a user-defined camera trajectory. PaintScene4D outperforms existing baselines in terms of visual quality, 3D consistency, and motion accuracy while also providing the capability for explicit camera control during inference. Notably, our approach is the first to implement text-to-4D scene generation using open-source models, offering a valuable contribution to the field.
@@ A Demo Video
We have provided a project webpage at https://paintscene4d.github.io/, including a video demonstration. The demo video presented in the website highlights the versatility and robustness of our framework by showcasing a diverse gallery of results generated from various text prompts and scene configurations. Furthermore, to substantiate the effectiveness of our approach, the demo video includes detailed comparisons with baseline methods such as 4D-fy [2] and Dream-in-4D [62]. Although both 4D-fy and Dream-in-4D are designed for object-level generation, they require significantly more computational time than our approach. Furthermore, these methods are limited to generating object-level renderings, whereas our framework is capable of rendering complete scene-level generations. These comparisons visually demonstrate the superior fidelity, consistency, and dynamicity achieved by our method across a wide range of scenarios.
In the demo video, we also showcase explicit control over camera movements in a rendered scene, offering a significant advantage over Text-to-Video (T2V) models. T2V models lack direct camera control, relying instead on implicit instructions through text prompts, which often yield inconsistent and less effective results. Additionally, T2V models generate a different scene for each iteration, even with identical prompts, limiting reproducibility. In contrast, our method supports precise manipulation of the camera trajectory within the same scene, ensuring consistency and offering greater flexibility for tailored visual outputs.
@@ B More Qualitative Results
In Figure B and Figure C, we provide more examples generated using our proposed framework to demonstrate the robustness of our methodology. The horizontal axis represents the time axis, and the vertical axis represents different view points. To fully appreciate the quality and diversity of our text-to-4D generation results, we strongly recommend viewing the accompanying video.
@@ C.1 Hyperparameters
Warping and Inpainting Module.   Table A demonstrates the parameters used in the warping and inpainting module. The values are carefully selected to balance efficiency and quality. The Number of Cameras determines the multi-view coverage necessary for generating high-quality reconstructions. The Relative Depth Estimator and Absolute Depth Estimator are key for warping operation, with DepthCrafter used for relative depth estimation and Metric3D v2 for absolute scaling. We inpaint the missing regions multiple times and pick the best one using a CLIP [35] based selector. The Inpainting Iterations represents the number of times we inpaint the missing region. The Filter Size for Bilateral Filtering sharpen the edges of a depth map necessitating better inpainting quality.
@@
Table B: Hyperparameters for 4D gaussian splatting training process.
@@
Table C: Hyperparameters for generating video from text-to-video model CogVideoX [54].
@@
Text-to-Video Generation.   Table C presents hyperparameters for video generation with the text-to-video model CogVideoX [54]. The hyperparameters are chosen to optimize the synthesis quality and temporal consistency. The Number of Timestamps defines the temporal resolution of the scene. The Number of Inference Steps impacts generation fidelity.
@@
Figure A: Failure Case: Our method is dependent on the assumption that the initial video generation exhibits no large camera movement. Large camera motion in the video, introduces distortions and artifacts during the subsequent rendering process, significantly affecting the visual fidelity of the final output.
Figure B: Gallery of Results: We present qualitative results of our text-to-4D generation framework, showcasing superior visual fidelity, consistent multi-view reconstructions, plausible scene compositions, and realistic dynamic motions. The horizontal axis represents the time axis and the vertical axis represents different view points. A comprehensive collection of video demonstrations is provided in the supplementary materials.
Figure C: Gallery of More Results. A comprehensive collection of video demonstrations is provided in the supplementary materials.
D Limitations and Future Work

### v2

@@ 3.3 Consistent Inpainting Module (CIM)
Figure 4: Comparisons with state-of-the-art text-to-4D generation methods. While both baseline methods produce scenes that broadly align with the text prompts, they lack essential fine details. Specifically, 4D-fy shows minimal motion and limited detail, whereas Dream-in-4D captures dynamics more effectively but produces stylized, cartoon-like renderings. In contrast, our method synthesizes photorealistic 4D scenes that faithfully follow the input text prompt while presenting significant, realistic dynamics within the scene.
Figure 5: Comparison against 4Real. [58] We demonstrate that our method produces more dynamics, larger scene coverage and better video-text alignment, and overall realism scenes.
Figure 6: Comparison against VividDream. [21] PaintScene4D demonstrates superior performance compared to VividDream in terms of higher quality (row 1), more dynamics (row 2), and higher motion and overall realism (row 3).
@@ 3.4 Training and Optimization
After performing all warping and inpainting operations across views and timestamps, we establish a comprehensive camera network, where each camera contains video frames captured from its respective viewpoint. Importantly, this multi-view setup is constructed without the need for model-specific training. Using this multi-view spatial information and temporal dynamics, we employ a 4D Gaussian rendering approach [51] to synthesize novel perspectives of the scene. The renderer takes Gaussian parameters, along with the timestamp, and computes the timestamp-conditioned deformation of these parameters. This approach enables continuous modeling of deformation, facilitating smooth interpolation between timestamps during novel view synthesis. At test time, any desired viewpoint and timestamp can be selected to generate a novel view.
@@ 4 Experiments
In this section, we introduce the baselines to compare with, the evaluation metrics, qualitative and quantitative results, and the analysis. The implementation details and additional analysis are presented in the supplementary material.
Figure 7: Camera Control. PaintScene4D shows strong explicit camera control capabilities. To guide T2V models (e.g., CogVideoX [55]), we append camera motion directives to the text prompt such as “The camera tilts to the right / upwards.” However, due to the implicit handling of camera motion, T2V often fails to generate precise or controllable camera movements. Our approach, once trained, allows for flexible camera trajectories within the bounds of the input cameras, achieving precise and repeatable control over camera movements.
Figure 8: Qualitative Results on real-world videos. Our work can produce realistic and coherent results with real-world video inputs, rather than limited to generated videos from T2V models.
@@ 4.1 Baselines and Evaluation Metrics
In the absence of open-source implementations for text-to-4D scene-level generation, we benchmark our approach against state-of-the-art text-to-4D object-level generation methods, namely 4Dfy [2] and Dream-in-4D [62], across a varied set of 20 prompts. We also compare against closed-source scene-level models [58, 21] with examples shown in their paper using the same text prompts. To assess the effectiveness of our proposed approach, we use the CLIP Score [35] alongside a structured user study. More visualization and comparisons with other closed-source models [44, 3] are included in the supplementary material.
CLIP Score.   The CLIP score [32] assesses the alignment between textual prompts and visual contents by calculating the cosine similarity between their embeddings [35]. Scored between 0 and 100, a higher value indicates a closer match. We compute CLIP scores by evaluating each frame with CLIP ViT-B/32 and averaging the scores across all frames and prompts for consistency.
User Study.   A comprehensive user study is conducted using Google Forms, involving 30 evaluators per video pair. Each evaluator receives three anonymized videos, each capturing a dynamic scene from a camera moving along a circular trajectory. The videos are accompanied by the original text prompt. The evaluators were asked to rate their preferences based on four criteria: motion realism, video-text alignment, high dynamicity, and general realism.
4.2 Text-to-4D Generation
Table 1: Quantitative results. We compare our method against object-level methods [2, 62] (above the dotted lines) and against scene-level methods [58] (below the dotted lines). The methods are evaluated in terms of CLIP score and human preference, including motion realism (MR), video-text alignment (VTA), high dynamicity (HR), general realism (GR), and overall preference. The reported human preference is the percentage of users who voted for the respective method in a head-to-head comparison.
@@ 4.2 Text-to-4D Generation
[…]
4D-fy [2] 31.8 2% 11% 5% 7% 7%
Dream-in-4D [62] 28.1 13% 14% 17% 2% 11%
PaintScene4D 36.0 85% 75% 78% 91% 82%
[…]
PaintScene4D 35.5 41% 58% 81% 61% 66%
Qualitative Results.   In Figure 4, we visualize spatio-temporal renderings produced by our method compared to baselines. Although all approaches are capable of synthesizing 4D scenes, baselines focus on object-level renderings and lack fine spatial details. Our approach, by contrast, generates scene-level 4D reconstructions in a significantly reduced time, producing realistic renderings. Notably, 4D-fy struggles to model realistic motion, while Dream-in-4D produces cartoonish effects that diminish realism.
We also compare our approach with closed-source scene generation models, 4Real and VividDream, as shown in Figure 5 and Figure 6. Our observations indicate that 4Real tends to generate outputs with a more cartoon-like appearance and limited scene dynamics. Additionally, its resolution is restricted due to constraints imposed by SDS-based optimization. Similarly, while VividDream produces more realistic results, it struggles to capture the level of dynamic motion typically observed in real-world scenarios. In contrast, our method achieves high photorealistic quality across both spatial and temporal dimensions. A gallery showcasing our results is presented in Figure 3.
Quantitative Results.   The CLIP score and the user study are reported in Table 1. In both object- and scene-level comparisons, our method outperforms all baseline models. The evaluations show a statistically significant preference for PaintScene4D due to its higher motion realism, photorealistic rendering of both the foreground and background, overall realism, and better video-text alignment.
Figure 9: Ablation Study. We demonstrate that each of our proposed components is essential for mitigating artifacts and inconsistencies, resulting in smooth consistent renderings.
Table 2: Comparison of computation efficiency. Modules time (hr) $\downarrow$ scene init 0.2 view warping 0.2 inpainting 1.0 4D rendering 0.8 PaintScene4D 2.2 4D-fy [2] 23 Dream4D [62] 4.5 4Real [58] 3.5 VividDream [21] 3.5
Runtime Analysis.   Table 2 presents a comparison of the time required to render a 4D scene from a given text prompt. Compared to object- and scene-level models, our method achieves the fastest training and rendering speed attributed to our modular and training-free framework. Additionally, to ensure a fair comparison, we normalize the rendering time based on a similar output resolution (720p).
@@ 4.3 Explicit Camera Control
To assess camera control, we compare our framework with other text-to-video (T2V) models, as illustrated in Figure 7. We input the same text prompts into the T2V model twice, adjusting only the camera movement description to direct it to “tilt towards the right” in one case and “move upwards” in the other. Our observations reveal two key limitations of the T2V models. First, even with a fixed seed, the T2V model generates different scenes for each altered prompt. Second, although the model simulates an upward camera movement in the second case, it lacks explicit control over the degree of camera motion. In contrast, our approach enables explicit, consistent control over camera trajectory within the same scene and motion dynamics, leveraging 4D modeling for precise camera manipulation.
4.4 Results on Real-World Videos
Our proposed method exhibits strong generalization capabilities when applied to real-world video data, as demonstrated in Figure 8. In contrast to approaches that are strictly limited by the output of text-to-video models, our method effectively processes real video input, making it more versatile and applicable to a wider range of scenarios. A key strength of our approach is its ability to extend beyond the frames of the initial video to construct a complete 4D scene. This means that it does not just reconstruct what is visible in the given video but infers and generates additional spatial-temporal information to create a more comprehensive and dynamic scene. This ability is crucial for applications requiring realistic and immersive 4D reconstruction.
@@ 5 Ablation Study
Table 3: Ablation Study. Model CLIP $\uparrow$ w/o CIM Module 30.8 w/o Farthest View 31.2 w/o Depth Alignment 33.9 PaintScene4D 36.0
Integrating video generation and inpainting modules is non-trivial and requires careful technical design to ensure high-quality results. Our ablation study (Figure 9 and Table 3) demonstrates that the removal of key components introduces significant artifacts and inconsistencies.
Depth Alignment Module.   The inclusion of the depth alignment module is crucial for maintaining the geometric consistency of the foreground. During the warping process, all frames are utilized, and any depth inconsistencies across frames result in error accumulation, leading to noticeable artifacts, particularly at the foreground boundaries.
Farthest View Sampling.   In PaintScene4D, we select the farthest view at each step of the warping process to maximize the inpainted area. Omitting this step causes severe degradation near the edges of the foreground, such as the panda’s boundary, where needle-shaped artifacts emerge due to the Gaussian splatting process.
Consistent Inpainting Module.   Temporal consistency in inpainting is essential for coherent 4D scene generation. Without this module, inpainting becomes inconsistent at the boundaries of objects (e.g., the panda) across different timestamps, leading to significantly degraded renderings.
@@ 6 Conclusion
We introduce PaintScene4D, a novel framework for generating photorealistic 4D scenes from a single text prompt. Our method addresses the challenges of spatial and temporal inconsistencies and enables the generation of novel views along a user-defined camera trajectory. PaintScene4D outperforms existing baselines in terms of visual quality, 3D consistency, motion accuracy, and generation efficiency. Comprehensive evaluations show leading results in the generation of 4D scenes, with significantly reduced computational requirements and enhanced camera control options.
@@ A Demo Video
We have provided a project webpage at https://paintscene4d.github.io/, including a video demonstration. This video highlights the versatility and robustness of our framework with a diverse gallery of results generated from various text prompts and scene configurations. Furthermore, to substantiate the effectiveness of our approach, the video includes detailed comparisons with baseline methods such as 4D-fy [2] and Dream-in-4D [62]. Although both 4D-fy and Dream-in-4D are designed for object-level generation, they require significantly more computational time than our approach. Furthermore, these methods are limited to generating object-level renderings, whereas our framework can render complete scene-level generations. These comparisons visually demonstrate the superior fidelity, consistency, and dynamicity achieved by our method across a wide range of scenarios.
In the video, we also showcase explicit control over camera movements in a rendered scene, offering a significant advantage over text-to-video (T2V) models. T2V models lack direct camera control, instead relying on implicit instructions through text prompts, which often yield inconsistent and less effective results. Additionally, T2V models generate a different scene for each iteration, even with identical prompts, limiting reproducibility. In contrast, our method supports precise manipulation of the camera trajectory within the same scene, ensuring consistency and offering greater flexibility for tailored visual outputs.
@@ B More Qualitative Results
In Figure C and Figure D, we provide more examples generated using our proposed framework to demonstrate the robustness of our methodology. The horizontal axis represents the time axis, and the vertical axis represents different viewpoints. To fully appreciate the quality and diversity of our text-to-4D generation results, we strongly recommend viewing the accompanying video.
@@ C PaintScene4D: Implementation Details
Our optimization framework comprises two stages: initially reconstructing a network of cameras, each associated with its respective view of the time frame, followed by training a 4D renderer. Specifically, we construct a network of 25 cameras and utilize videos that span 50 timestamps. All experiments are performed on a single A100 GPU. The complete warping and inpainting process, which is performed without additional training, requires approximately two hours. Following this, the 4D renderer is trained in about one hour, resulting in a total of approximately 3 hours to complete the training and generate novel views along any desired trajectory. This duration is significantly shorter than the training time required by recent state-of-the-art methods: Dream-in-4D takes over four hours, while 4Dfy takes over 20 hours, despite producing only object-level 4D renderings. To initialize the scene and establish motion priors for 4D reconstruction, we use CogVideoX-5b [55]. For depth estimation, DepthCrafter [16] is used, as it produces consistent depth estimates across video frames, enabling reliable warping. Perspective Fields [20] is used to estimate the camera intrinsics for the generated video. For the segmentation model to distinguish foreground and background, we use GroundingSAM-2 [37].
@@ C.1 Hyperparameters
Warping and Inpainting Module.   Table A demonstrates the parameters used in the warping and inpainting module. The values are carefully selected to balance efficiency and quality. The Number of Cameras determines the multi-view coverage necessary for generating high-quality reconstructions. The Relative Depth Estimator and Absolute Depth Estimator are key for warping operation, with DepthCrafter used for relative depth estimation and Metric3D v2 for absolute scaling. We inpaint the missing regions multiple times and pick the best one using a CLIP [35] based selector. The Inpainting Iterations represents the number of times we inpaint the missing region. The Filter Size for Bilateral Filtering sharpens the edges of a depth map necessitating better inpainting quality.
@@
Table B: Hyperparameters for 4D-GS rendering process.
@@
Table C: Hyperparameters for CogVideoX [55] video generation.
@@
Text-to-Video Generation.   Table C presents hyperparameters for video generation with the CogVideoX text-to-video model [55]. The hyperparameters are chosen to optimize the synthesis quality and temporal consistency. The Number of Timestamps defines the temporal resolution of the scene. The Number of Inference Steps impacts generation fidelity.
D More Comparison with Other Methods
Additionally, we compare our method with other object-level approaches, such as MAV3D and TC4D, as shown in Figure A. These methods are limited to generating renderings at the object level, restricting their ability to capture broader scene contexts. Our results demonstrate that, in both cases, our method produces renderings with higher realism while also achieving faster generation times. This efficiency is notable given that MAV3D and TC4D are constrained to object-level synthesis, whereas our approach extends beyond these limitations.
E Discussion on Success Rate
Our method demonstrates robustness to minor camera movements in the input video, ensuring stability in most cases. However, when there are significant deviations from a static camera setup, failure cases may occur, as illustrated in Supplementary Figure B. To maintain a static camera view, an appropriate prompt design can be utilized, as discussed in Section 4.1 of the main paper. With a success rate exceeding 90%, we present our results without selective curation, further emphasizing the reliability and consistency of our approach in generating high-quality 4D scenes.
Figure A: Comparison with additional text-to-4D methods.
@@ E Discussion on Success Rate
Figure B: Failure Case: Our method is dependent on the assumption that the initial video generation exhibits no large camera movement. Large camera motion in the video, introduces distortions and artifacts during the subsequent rendering process, significantly affecting the visual fidelity of the final output.
F Limitations and Future Work
While our method successfully generates photorealistic 4D scenes from a single text prompt, several limitations persist:
  1. 1.